Please verify that sin(x) + tan(x)/ 1 + sec(x) = sin(x)
1
Expert's answer
2012-03-25T09:25:51-0400
(sin(x) + tan(x))/( 1 + sec(x)) = sin(x)
(sin(x) +sin(x)/cos(x))/( 1 + 1/cos(x)) = sin(x)

(sin(x)cos(x) +sin(x))/cos(x) /(( cos(x) + 1)/cos(x)) = sin(x)
sin(x)(cos(x) +1)/ /( cos(x) + 1) = sin(x)

sin(x)=sin(x)